OpenStudy (shubhamsrg):

how do i prove n^4 + 4^n is never prime for n>1

OpenStudy (anonymous):

Proof by induction? Show the statement holds for \(n=2\), then (assuming it's true for \(n=k\)) show that it holds for \(n=k+1\).

OpenStudy (shubhamsrg):

well yes I tried and am stuck there, where we prove for n=k+1

OpenStudy (shubhamsrg):

i have even tried this way, clearly its true when n=2m for some natural no. m but again stuck where we need to prove for n=2m+1

OpenStudy (anonymous):

For the \(n=k+1\) case, you have \((k+1)^4+4^{k+1}\). Try expanding the binomial and see what happens for even and odd \(k\).

OpenStudy (shubhamsrg):

well i have assumed 4^k + k^4 = ab where a,b>1 now for k+1 part, 4^(k+1) + (k+1)^4 => 3.4^k + ab + 4k^3 + 6k^2 + 4k +1 stuck here :/

OpenStudy (anonymous):

Do you know about modular arithmetic?

OpenStudy (shubhamsrg):

bit

OpenStudy (anonymous):

Clearly if n is even, the result will be divisible by 2. So you should only worry about showing that its not prime if n is odd. Modular arithmetic can take care of that case.

OpenStudy (shubhamsrg):

please guide me through ? not completely, but hint me please

OpenStudy (anonymous):

Try plugging in some odd numbers. Try to guess what always divides to the statement if n is odd.

OpenStudy (shubhamsrg):

well an interesting result that comes is that this f(n) is always multiple of 2 or more different primes

OpenStudy (anonymous):

First note that every integer \(n>1\) is either even or odd. If \(n\) is even, then \(n=2k\) for some positive integer \(k\), and thus \(n^4+4^n=2^4k^4+4^{2k}=2(2^3k^4+2^{4k-1})\), so clearly \(2\mid n^4+4^n\) and therefore \(n^4+4^n\) is not prime. If \(n\) is odd, then \(n=2k+1\) for some positive integer \(k\). Consider \(4^n\mod5\) and it should be trivial.

OpenStudy (shubhamsrg):

what do we do about the polynomial ?

OpenStudy (shubhamsrg):

and yes i see by hit and trial that f(n) is always divisible by 5 !! :O

OpenStudy (anonymous):

Euler's theorem tells us that \(4^n=4^{n\mod4}\mod5\). Consider, then:$$4^0=1\mod5\\4^1=4\mod5\\4^2=1\mod5\\4^3=4\mod 5$$Since \(n\) is odd, \(n=1,3\mod4\) and thus \(4^n=4\mod5\). This leaves us to show that \(n^4=1\mod5\). Recognize $$n^4=(2k+1)^4=(2k)^4+4(2k)^3+6(2k)^8+4(2k)+1=1\mod5$$... and we're done.

OpenStudy (shubhamsrg):

not too sure how you directly conclude (2k)^4 +4(2k)^3 + 6(2k)^2 + 4(2k) is divisible by 5 ?

OpenStudy (anonymous):

Can you show $$n^4-1=(n^2+1)(n^2-1)=(n^2+1)(n-1)(n+1)=0\mod5$$?

OpenStudy (anonymous):

(2k)^2 is what I meant, btw.

OpenStudy (anonymous):

try this also: for even numbers its true clearly...for odd numbers\[n^4+4^n=n^4+4^{2k+1}=n^4+4\times 4^{2k}=n^4+4\times (2^k)^4=n^4+4m^4\]try to factor last expression

OpenStudy (shubhamsrg):

oh of course! (n^2 -2mn + 2m^2 ) (n^2 +2mn + 2m^2) hence not prime !! thank you @mukushla and am sorry i am still unable to follow that mod5 method @oldrin.bataku :(

OpenStudy (anonymous):

$$n+1=0\mod5\implies n=4\mod5\\n-1=0\mod5\implies n=1\mod 5\\n^2+1=0\mod5\implies n^2=4\mod5\implies n=2,3\mod 5$$So we've established $$n=1,3\implies n^4-1=0\mod5$$and thus $$n^4=1\mod 5$$

OpenStudy (shubhamsrg):

confused at n^2 + 1 = 0 part ?

OpenStudy (shubhamsrg):

oh no got it

OpenStudy (anonymous):

We factored \(n^4-1=(n^2+1)(n-1)(n+1)\) and showed that no matter what \(n\) is we end up with \(n^4-1=0\mod5\).

OpenStudy (shubhamsrg):

except for n=5m ?